top of page

Depot Report - Saturday 20th January 2024


55009

Work continued on circuit testing on the electrics with more to do next weekend. PU 408 was started and ran Ok.

One of the engine room door catches was repaired so all  now close securely. 



55015 and 55019

Both due their ultrasonic axle & equalising beam testing so 2 teams of 2, 1 on each loco, set about removing the axle box covers and the inner plate and cleaned out all of the old grease that was accessible prior to the new grease being put in. The covers were then replaced to keep dirt out until the inspection date. New tab washers arrived and gaskets on the way. No other work on 19.



55015.

2 floor panels fitted in No 1 cab, Nose End door lock in No 2 end now works so that's all internal door locks now working. New piping was fitted in No 2 end for the new screen wash units. One of the collector drum drain plate cover was removed along with another on a spare from the stores so that drain taps can be fitted and the system can then drain into the bedplate. The last of the traction motor cable junction boxes was fitted to No 6 motor.

Thanks to Simon, Andrea. Geoff, Nigel, Richard, Peter, Bill, Steve, Paul, Richie, Daniel.

Photos: Mike Hallam-Rudd



252 views0 comments

Recent Posts

See All

Comments


bottom of page